OpenStudy (dtan5457):

Help finalizing solving tan 195 degrees

OpenStudy (dtan5457):

Using the difference of tan i put tan 240-tan 45

OpenStudy (dtan5457):

that's basically tan 60-tan 45 tan60=sqrt of 3 tan45=1

OpenStudy (dtan5457):

i'm stuck at \[\frac{ -4+2\sqrt{3} }{ -2 }\]

OpenStudy (dtan5457):

oh wait would this be \[2-\sqrt{3}\]

OpenStudy (anonymous):

yes

OpenStudy (dtan5457):

ty

OpenStudy (greencat):

 The reference angle is 60, but that does not always mean that you can replace it with 60. In this case, you could replace tan 240 with tan 60 because of them both being positive (All Silver Tea Cups). In the future, you should take caution.